Incident angle is 30 and refracted angle is 45. What is the refractive index of denser medium with respect to rarer medium?

Grade:12th pass

1 Answers

Mahima Kanawat
5 years ago
Dear student 
This can be find out by Snell's law
Take refractive index of rarer medium as 1 ( just to compare both media)
n1sin i = nsin r
1 * sin 30° = n2 sin 45°
1*1/2 = n* 1/ (2)1/2
n= 1 / (2)1/2
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT